# An eclipsing 8.56 minute orbital period mass-transferring binary

### v2

@@ Abstract
We report the discovery of ATLAS J101342.5-451656.8 (hereafter referred to as ATLAS J1013-4516), an 8.56 minute orbital-period mass-transferring AM Canum Venaticorum (AM CVn) binary with a mean Gaia magnitude of $G=19.51$, identified via periodic variability in light curves from the Asteroid Terrestrial-impact Last Alert System (ATLAS) of Gaia white-dwarf candidates. Follow-up with the Large Lenslet Array Magellan Spectrograph (LLAMAS) shows a helium-dominated accretion disk, and high-speed ULTRACAM photometry reveals pronounced primary and secondary eclipses. We construct a decade-long timing baseline leveraging lightcurves from the ATLAS and Gaia surveys, as well as the high-speed imagers ULTRACAM on the NTT and proto-Lightspeed on the Magellan Clay telescope. From this timing baseline, we measure an orbital period derivative of $\dot{P}=-1.60\pm 0.07\times 10^{-12}$ s s-1. Interpreted in the context of stable mass transfer, the magnitude and sign of $\dot{P}$ indicate that the orbital evolution is governed by the interplay between gravitational-wave–driven angular-momentum losses and mass transfer, directly probing the donor’s structural response to mass loss. We constrain the accretor and donor mass based on stable mass-transfer arguments assuming angular momentum loss dominated by gravitational wave emission, allowing us to infer the characteristic gravitational wave strain of the binary for future space-based GW observatories such as the Laser Interferometer Space Antenna (LISA). We predict a characteristic strain corresponding to a 4-year LISA signal-to-noise ratio $\gtrsim 20$, establishing ATLAS J1013-4516 as a strong prospective LISA source that will probe long-term orbital evolution in the mass-transferring regime.
@@ I Introduction
To date, only $\sim$100 AM CVn binaries have been identified, and the majority have orbital periods $\gtrsim 20$ min (Solheim, 2010; Green et al., 2025). The known sample has been assembled primarily through a combination of spectroscopy (identifying blue continua with strong He I/He II features and no Balmer lines), the detection of dwarf nova outbursts in transient surveys, and searches for periodic variability in time-domain data, and X-ray–selected ultracompact binaries (e.g., Israel et al., 1999; Solheim, 2010; Burdge et al., 2020; Green et al., 2025). The handful of AM CVn binaries with periods below $\sim 10$ min lie where gravitational-wave–driven angular-momentum loss is strongest, making them some of the highest–strain-amplitude Galactic sources in the millihertz band. Many are expected to be individually resolved by LISA, and their numbers and period distribution inform models of the unresolved Galactic gravitational-wave foreground (Korol et al., 2017; Kupfer et al., 2018). These ultracompact, mass-transferring binaries also probe helium mass-transfer rates and stability in double-degenerate binaries, providing empirical constraints on the conditions that separate stable accretion from the regimes that produce helium-powered thermonuclear transients (Webbink, 1984; Bildsten et al., 2007; Shen & Bildsten, 2009; Piersanti et al., 2014; Munday et al., 2025). The advent of wide-field, high-cadence surveys—including the Asteroid Terrestrial-impact Last Alert System (ATLAS; Tonry et al. (2018)), the Zwicky Transient Facility (ZTF; Bellm et al. (2019)), the Gaia mission (Gaia Collaboration et al., 2016), and the Transiting Exoplanet Survey Satellite (TESS; Ricker et al. (2015))—now enables systematic searches for such short-period sources across the sky (Burdge et al., 2020; Green et al., 2025).
@@ II.2 Gaia Epoch Photometry
In addition to ATLAS, we make use of time-resolved photometry from the Gaia mission (Gaia Collaboration et al., 2016). We retrieve Gaia G-band epoch photometry from Data Release 3 (2023A&A...674A...1G), which provides 54 sparsely sampled but high-precision flux measurements spanning $\sim$2014–2017. The astrometric parallax of ATLAS J1013-4516 is poorly constrained ($0.1992\pm 0.2808$ mas) and is not used to infer the system distance.
@@ III.1 Optical Spectral Features
The coadded Magellan/LLAMAS spectrum of ATLAS J1013-4516 (Figure 2) is dominated by strong He ii emission, most prominently at 4542, 4686, 5411, and 6527 Å. The phase-resolved line profiles (Figure 2, right panels) exhibit persistent double-peaked morphologies consistent with Doppler broadening in a rotating accretion disk. Mild asymmetries in the relative peak strengths are visible at some orbital phases (e.g., $\phi=0.0$ and $\phi=0.5$ in the He ii 4686 Åprofile), indicating departures from a perfectly axisymmetric brightness distribution. However, within the signal-to-noise and spectral resolution of the LLAMAS data, we do not detect a coherent phase-dependent velocity modulation or a distinctive S-wave component. This implies that any localized emission regions (e.g., a stream–disk impact or azimuthal disk structure) are weak, spatially extended, or diluted by the dominant disk emission. No Balmer emission or absorption features are detected, confirming a helium-dominated accretion flow.
Several weak features are present near the expected wavelengths of He i transitions (e.g., 4471, 5015, 5876, and 6678 Å), but their low amplitudes and limited signal-to-noise preclude a reobust identification. The relative prominence of He ii over He i suggests a hot, high-excitation environment, as seen in the most luminous AM CVn systems such as ES Cet (Espaillat et al., 2005) and ZTF J0546+3843 (Chakraborty et al., 2024).
Weak emission near $\sim$3995 Å and the 4630–4640 Å region may correspond to N ii transitions, though these features are detected at marginal significance. No statistically significant C or O emission lines are observed. Higher signal-to-noise spectroscopy will be required to place meaningful constraints on the elemental abundances. We now turn to the time-domain data, which reveal the system’s orbital geometry and reveal non-axisymmetric emission components that are not cleanly separable in the phase-averaged spectrum.
@@ III.2 Light Curve Morphology and Modeling
The post-eclipse hump in ATLAS J1013-4516 may also signal contributions from a vertically extended multi-site emission associated with the stream–disk interaction. In hydrogen-accreting cataclysmic variables, partial stream overflow—where material skims the outer disk rim and later re-impacts the disk at smaller radii—can produce secondary emission regions displaced in azimuth from the primary hot spot (Hessman, 1999; Speith & Kunze, 2002). Eclipse mapping of novalike systems such as UU Aqr and IY UMa has revealed enhanced emission along the ballistic stream beyond the initial impact site, consistent with such disk-skimming components (Baptista et al., 2000). In the helium-dominated ultracompact Gaia14aae, phase-resolve spectroscopy has revealed two bright spots, one consistent with the canonical stream–disk impact and a second likely associated with re-impact or vertically extended overflow (Green et al., 2019). Comparable behavior has also been reported in other short-period systems such as SDSS J0926++3624 (Schlindwein & Baptista, 2018). A modest contribution from analogous structures could contribute to the observed out-of-eclipse asymmetry in ATLAS J1013-4516, although the present data do not uniquely require this interpretation. Alternative mechanisms—including magnetic channeling, disk warping, or anisotropic irradiation—-could also produce phase-dependent asymmetries in disk emission. Given the degeneracy among these scenarios, the observed phase offsets primarily constrain the azimuthal location of enhanced emission rather than its detailed physical origin.
@@ IV.2 Orbital Period Derivative and Donor Physics
In Roche-lobe–filling binaries, orbital angular momentum losses (e.g., gravitational radiation) act concurrently with mass transfer, which alters the component masses and thereby modifies the orbital response to a given angular-momentum loss. As the donor loses mass, its short-timescale structural response—quantified by the adiabatic mass-radius exponent $\xi_{\rm ad}\equiv\rm d\ln R_{2}/\rm d\ln M_{2}$—determines whether the donor expands or contracts, while the evolving mass ratio alters the Roche-lobe radius. Continued Roche-lobe overflow therefore couples angular momentum loss to mass transfer, such that for degenerate donors ($\xi_{\rm ad}<0$) the orbit can widen even as the total orbital angular momentum decreases. This coupling gives rise to the characteristic period minimum in AM CVn systems, separating an early phase of orbital decay from a later phase of orbital expansion once the donor’s adiabatic response becomes sufficiently negative (e.g., Nelemans et al. 2001; Deloye & Bildsten 2003).
@@
To assess the physical plausibility of the inferred parameter space, we repeat the analysis using finite-entropy helium white dwarf donor models, in which $(M_{2},R_{2},\xi_{\rm ad})$ are linked self-consistently through the finite-temperature relations of Deloye & Bildsten (2003) and required to satisfy Roche-lobe contact. Under this assumption, the allowed ranges of $\xi_{\rm ad}$ and $\chi_{\rm evol}$ collapse to a comparatively narrow region of parameter space (Figure 9), indicating that if the donor is a warm helium white dwarf, the observed orbital-period derivative meaningfully constrains its effective radius response to mass loss. The inferred component masses remain broadly consistent with the EOS-agnostic and finite-entropy analyses, suggesting that the mass constraints are robust to uncertainties in the donor’s thermal state; we therefore adopt the finite-entropy helium white dwarf model and summarize the corresponding marginalized posterior constraints in the Table 2.
@@
Figure 9: Comparison of mass and donor-physics constraints for ATLAS J1013-4516 under different assumptions about the donor structure. Top: EOS-agnostic analysis in which the donor’s adiabatic mass–radius exponent $\xi_{\rm ad}$ and evolutionary ratio $\chi_{\rm evol}$ are treated as free parameters. Bottom: Finite-entropy helium white dwarf donor model based on the relations of Deloye & Bildsten (2003), which self-consistently link $(M_{2},R_{2},\xi_{\rm ad})$ and enforce Roche-lobe contact. Contours indicate 68% and 95% credible regions, while hatched areas denote excluded parameter space. While the inferred component masses are broadly consistent across both treatments, adopting a physically motivated donor equation of state substantially tightens the allowed ranges of $\xi_{\rm ad}$ and $\chi_{\rm evol}$, reflecting the reduced freedom in the donor’s mass–radius evolution.
@@ IV.3 Population Context and Comparisons
This apparent discrepancy may be reconciled if the observed optical flux of ATLAS J1013-4516 is strongly suppressed by inclination-dependent obscuration or reprocessing. High-inclination accreting systems are known to exhibit reduced observed luminosities due to vertically extended disk rims, stream–disk impact regions, or disk winds that obscure or redistribute accretion power out of the optical bandpasses, an effect well documented in high-inclination X-ray binaries (e.g., White & Holt 1982). Consistent with this interpretation, ES Cet exhibits a partial eclipse of the outer accretion disk, indicative of a lower inclination, whereas the deep, sharp eclipses observed in ATLAS J1013-4516 require a more edge-on geometry. A higher inclination would naturally enhance self-obscuration, allowing a system with a high intrinsic accretion luminosity to appear optically faint.

### v3

@@ Abstract
We report the discovery of ATLAS J101342.5-451656.8 (hereafter referred to as ATLAS J1013-4516), an 8.56 minute orbital-period mass-transferring AM Canum Venaticorum (AM CVn) binary with a mean Gaia magnitude of $G=19.51$, identified via periodic variability in light curves from the Asteroid Terrestrial-impact Last Alert System (ATLAS) of Gaia white-dwarf candidates. Follow-up with the Large Lenslet Array Magellan Spectrograph (LLAMAS) shows a helium-dominated accretion disk, and high-speed ULTRACAM photometry reveals pronounced primary and secondary eclipses. We construct a decade-long timing baseline leveraging lightcurves from the ATLAS and Gaia surveys, as well as the high-speed imagers ULTRACAM on the NTT and proto-Lightspeed on the Magellan Clay telescope. From this timing baseline, we measure an orbital period derivative of $\dot{P}=-1.60\pm 0.07\times 10^{-12}$ s s-1. Interpreted in the context of stable mass transfer, the magnitude and sign of $\dot{P}$ indicate that the orbital evolution is governed by the interplay between gravitational-wave–driven angular-momentum losses and mass transfer, directly probing the donor’s structural response to mass loss. We constrain the accretor and donor mass based on stable mass-transfer arguments assuming angular momentum loss dominated by gravitational wave emission, allowing us to infer the characteristic gravitational wave strain of the binary for future space-based GW observatories such as the Laser Interferometer Space Antenna (LISA). We predict a characteristic strain corresponding to a 4-year LISA signal-to-noise ratio $\gtrsim 10$, establishing ATLAS J1013-4516 as a strong prospective LISA source that will probe long-term orbital evolution in the mass-transferring regime.
@@ I Introduction
To date, only $\sim$100 AM CVn binaries have been identified, and the majority have orbital periods $\gtrsim 20$ min (Solheim, 2010; Green et al., 2025). The known sample has been assembled primarily through a combination of spectroscopy (identifying blue continua with strong He I/He II features and no Balmer lines), the detection of dwarf nova outbursts in transient surveys, and searches for periodic variability in time-domain data, and X-ray–selected ultracompact binaries (e.g., Israel et al., 1999; Solheim, 2010; Burdge et al., 2020; Green et al., 2025). The handful of AM CVn binaries with periods below $\sim 10$ min lie where gravitational-wave–driven angular-momentum loss is strongest, making them some of the highest–strain-amplitude Galactic sources in the millihertz band. Many are expected to be individually resolved by LISA, and their numbers and period distribution inform models of the unresolved Galactic gravitational-wave foreground (Korol et al., 2017; Kupfer et al., 2018). These ultracompact, mass-transferring binaries also probe helium mass-transfer rates and stability in double-degenerate binaries, providing empirical constraints on the conditions that separate stable accretion from the regimes that produce helium-powered thermonuclear transients (Webbink, 1984; Bildsten et al., 2007; Shen and Bildsten, 2009; Piersanti et al., 2014; Munday et al., 2025). The advent of wide-field, high-cadence surveys—including the Asteroid Terrestrial-impact Last Alert System (ATLAS; Tonry et al. (2018)), the Zwicky Transient Facility (ZTF; Bellm et al. (2019)), the Gaia mission (Gaia Collaboration et al., 2016), and the Transiting Exoplanet Survey Satellite (TESS; Ricker et al. (2015))—now enables systematic searches for such short-period sources across the sky (Burdge et al., 2020; Green et al., 2025).
@@ II.2 Gaia Epoch Photometry
In addition to ATLAS, we make use of time-resolved photometry from the Gaia mission (Gaia Collaboration et al., 2016). We retrieve Gaia G-band epoch photometry from Data Release 3 (Gaia Collaboration et al., 2023), which provides 54 sparsely sampled but high-precision flux measurements spanning $\sim$2014–2017. The astrometric parallax of ATLAS J1013-4516 is poorly constrained ($0.1992\pm 0.2808$ mas) and is not used to infer the system distance.
@@ III.1 Optical Spectral Features
[…]
Several weak features are present near the expected wavelengths of He i transitions (e.g., 4471, 5015, 5876, and 6678 Å), but their low amplitudes and limited signal-to-noise preclude a robust identification. The relative prominence of He ii over He i suggests a hot, high-excitation environment, as seen in the most luminous AM CVn systems such as ES Cet (Espaillat et al., 2005) and ZTF J0546+3843 (Chakraborty et al., 2024).
Weak emission near $\sim$3995 Å and the 4630–4640 Å region may correspond to N ii transitions, though these features are detected at marginal significance. No statistically significant C or O emission lines are observed. Higher signal-to-noise spectroscopy will be required to place meaningful constraints on the elemental abundances. We now turn to the time-domain data, which constrain the system’s orbital geometry and reveal non-axisymmetric emission components that are not cleanly separable in the phase-averaged spectrum.
@@ III.2 Light Curve Morphology and Modeling
The post-eclipse hump in ATLAS J1013-4516 may also signal contributions from a vertically extended multi-site emission associated with the stream–disk interaction. In hydrogen-accreting cataclysmic variables, partial stream overflow—where material skims the outer disk rim and later re-impacts the disk at smaller radii—can produce secondary emission regions displaced in azimuth from the primary hot spot (Hessman, 1999; Speith and Kunze, 2002). Eclipse mapping of novalike systems such as UU Aqr and IY UMa has revealed enhanced emission along the ballistic stream beyond the initial impact site, consistent with such disk-skimming components (Baptista et al., 2000). In the helium-dominated ultracompact Gaia14aae, phase-resolve spectroscopy has revealed two bright spots, one consistent with the canonical stream–disk impact and a second likely associated with re-impact or vertically extended overflow (Green et al., 2019). Comparable behavior has also been reported in other short-period systems such as SDSS J0926++3624 (Schlindwein and Baptista, 2018). A modest contribution from analogous structures could contribute to the observed out-of-eclipse asymmetry in ATLAS J1013-4516, although the present data do not uniquely require this interpretation. Alternative mechanisms—including magnetic channeling, disk warping, or anisotropic irradiation—-could also produce phase-dependent asymmetries in disk emission. Given the degeneracy among these scenarios, the observed phase offsets primarily constrain the azimuthal location of enhanced emission rather than its detailed physical origin.
@@ IV.2 Orbital Period Derivative and Donor Physics
In Roche-lobe–filling binaries, orbital angular momentum losses (e.g., gravitational radiation) act concurrently with mass transfer, which alters the component masses and thereby modifies the orbital response to a given angular-momentum loss. As the donor loses mass, its short-timescale structural response—quantified by the adiabatic mass-radius exponent $\xi_{\rm ad}\equiv\rm d\ln R_{2}/\rm d\ln M_{2}$—determines whether the donor expands or contracts, while the evolving mass ratio alters the Roche-lobe radius. Continued Roche-lobe overflow therefore couples angular momentum loss to mass transfer, such that for degenerate donors ($\xi_{\rm ad}<0$) the orbit can widen even as the total orbital angular momentum decreases. This coupling gives rise to the characteristic period minimum in AM CVn systems, separating an early phase of orbital decay from a later phase of orbital expansion once the donor’s adiabatic response becomes sufficiently negative (e.g., Nelemans et al. 2001; Deloye and Bildsten 2003).
@@
To assess the physical plausibility of the inferred parameter space, we repeat the analysis using finite-entropy helium white dwarf donor models, in which $(M_{2},R_{2},\xi_{\rm ad})$ are linked self-consistently through the finite-temperature relations of Deloye and Bildsten (2003) and required to satisfy Roche-lobe contact. Under this assumption, the allowed ranges of $\xi_{\rm ad}$ and $\chi_{\rm evol}$ collapse to a comparatively narrow region of parameter space (Figure 9), indicating that if the donor is a warm helium white dwarf, the observed orbital-period derivative meaningfully constrains its effective radius response to mass loss. The inferred component masses remain broadly consistent with the EOS-agnostic and finite-entropy analyses, suggesting that the mass constraints are robust to uncertainties in the donor’s thermal state; we therefore adopt the finite-entropy helium white dwarf model and summarize the corresponding marginalized posterior constraints in the Table 2.
@@
Figure 9: Comparison of mass and donor-physics constraints for ATLAS J1013-4516 under different assumptions about the donor structure. Top: EOS-agnostic analysis in which the donor’s adiabatic mass–radius exponent $\xi_{\rm ad}$ and evolutionary ratio $\chi_{\rm evol}$ are treated as free parameters. Bottom: Finite-entropy helium white dwarf donor model based on the relations of Deloye and Bildsten (2003), which self-consistently link $(M_{2},R_{2},\xi_{\rm ad})$ and enforce Roche-lobe contact. Contours indicate 68% and 95% credible regions, while hatched areas denote excluded parameter space. While the inferred component masses are broadly consistent across both treatments, adopting a physically motivated donor equation of state substantially tightens the allowed ranges of $\xi_{\rm ad}$ and $\chi_{\rm evol}$, reflecting the reduced freedom in the donor’s mass–radius evolution.
@@ IV.3 Population Context and Comparisons
This apparent discrepancy may be reconciled if the observed optical flux of ATLAS J1013-4516 is strongly suppressed by inclination-dependent obscuration or reprocessing. High-inclination accreting systems are known to exhibit reduced observed luminosities due to vertically extended disk rims, stream–disk impact regions, or disk winds that obscure or redistribute accretion power out of the optical bandpasses, an effect well documented in high-inclination X-ray binaries (e.g., White and Holt 1982). Consistent with this interpretation, ES Cet exhibits a partial eclipse of the outer accretion disk, indicative of a lower inclination, whereas the deep, sharp eclipses observed in ATLAS J1013-4516 require a more edge-on geometry. A higher inclination would naturally enhance self-obscuration, allowing a system with a high intrinsic accretion luminosity to appear optically faint.